# On the isotopies of tangles in periodic 3-manifolds using finite covers

### v1

@@ 2.1. Links in 3-manifolds, periodic tangles, and minimal motifs
In [13], another definition of a minimal motif is given in terms of the associated periodic lattice. The existence and uniqueness of a minimal motif are non-trivial questions, which we will discuss in Section 5. Note that seeking systematic ways to find a minimal motif is still an open problem.
@@ 2.3. Prime decomposition
The connected sum $M_{0}\#M_{1}$ of 3-manifolds $M_{0}$ and $M_{1}$ is obtained by removing a ball in each $M_{i}$ and gluing them along the resulting boundary spheres. If $M_{0}$ and $M_{1}$ are connected and oriented, the oriented connected sum $M_{0}\#M_{1}$ is uniquely determined. The connected sum operation is commutative and associative, and $S^{3}$ is the identity element for the connected sum. A 3-manifold $M$ is prime if $M=P\#Q$ implies that $P=S^{3}$ or $Q=S^{3}$. A 3-manifold $M$ is irreducible if every embedded sphere in $M$ bounds a 3-ball. In the context of our paper, the 3-manifolds $S^{3}$, $S^{1}\times D^{2}$, $T^{2}\times I$, and $T^{3}$ are irreducible. If an orientable 3-manifold $M$ is prime and not irreducible, then $M=S^{1}\times S^{2}$. The prime decomposition of a compact orientable 3-manifold $M$ is a decomposition of the form $M=P_{1}\#\dots\#P_{n}$, where each $P_{i}$ is a prime 3-manifold. Moreover, a prime decomposition is unique up to permutation and connected sums with additional copies of $S^{3}$.
Let $L$ be a link in an irreducible 3-manifold $X$. We say that $L$ splits to links $L^{\prime}$, $L_{1}$, …, $L_{m}$ if $L=L^{\prime}\sqcup L_{1}\sqcup\dots\sqcup L_{m}$ and there are disjoint balls $B_{1},\dots,B_{m}$ in $X$ disjoint from $L^{\prime}$ such that each $B_{i}$ contains $L_{i}$. If $m\geq 1$, then $L$ is called a split link. The link $L$ is split if and only if the complement $X\setminus\!\!\setminus L$ is not irreducible. The following lemma is a consequence of the prime decomposition of the complement of link $L$, $X\setminus\!\!\setminus L$.
@@ 2.4. Seifert fibered 3-manifolds
For coprime integers $p$ and $q\neq 0$, a $(p,q)$-model Seifert fibration on $S^{1}\times D^{2}$ is a decomposition of $S^{1}\times D^{2}$ into disjoint circles, called fibers, obtained from the segments $[0,1]\times\{x\}$ in $[0,1]\times D^{2}$ by gluing $\{0\}\times D^{2}$ and $\{1\}\times D^{2}$ via a rotation of $2\pi p/q$. The $(p,1)$-model Seifert fibration is the same as the fibration given by the usual product structure on $S^{1}\times D^{2}$ (sometimes referred to as a product fibration) – that is, every fiber is a circle over a point in the disk and the resulting space is $S^{1}\times D^{2}$.
@@
The classification of Seifert fibered 3-manifolds was established many years ago, and here, we use the notation for orientable Seifert fibered 3-manifolds following [6, 25]. For $g\in\mathbb{Z}$ and $b\in\mathbb{Z}_{\geq 0}$, let $\Sigma_{g,b}$ denote the orientable compact surface of genus $g$ with $b$ boundary components if $g\geq 0$. Let $\Sigma_{g,b}$ denote the non-orientable compact surface that is the connected sum of $-g$ copies of $\mathbb{R}P^{2}$ and $\Sigma_{0,b}$ if $g<0$. Let $M(g,b;\alpha_{1}/\beta_{1},\dots,\alpha_{k}/\beta_{k})$ denote the orientable Seifert fibered 3-manifold that is fibered over $\Sigma_{g,b}$ with at most $k$ singular fibers and fiber data $\alpha_{i}/\beta_{i}\in\mathbb{Q}$. More precisely, the 3-manifold $M(g,b;\alpha_{1}/\beta_{1},\dots,\alpha_{k}/\beta_{k})$ is obtained from the orientable $S^{1}$-bundle over $\Sigma_{g,b+k}$ by Dehn filling along $k$ boundary components so that the meridians of attached solid tori are glued to the slopes $\alpha_{i}/\beta_{i}$. Here the slope $\alpha_{i}/\beta_{i}$ on the torus $T_{i}$ for coprime integers $\alpha_{i}$ and $\beta_{i}$ is represented by $\alpha_{i}\lambda_{i}+\beta_{i}\mu_{i}\in H_{1}(T_{i},\mathbb{Z})$ for an oriented pair of a fiber $\lambda_{i}$ and a boundary component of a section $\mu_{i}$. The Seifert fibration on $M(g,b;\alpha_{1}/\beta_{1},\dots,\alpha_{k}/\beta_{k})$ is obtained by extending the product fibration over $\Sigma_{g,b+k}$. The 3-manifolds with non-unique Seifert fibrations are classified. We state the classification for bounded orientable Seifert fibered 3-manifolds.
@@
The solid torus $S^{1}\times D^{2}$ admits the model Seifert fibrations of $M(0,1;\alpha/\beta)$. The thickened torus $T^{2}\times I$ admits the product fibration of $M(0,2;)$, which is unique up to isomorphism. The twisted $I$-bundle over the Klein bottle $S^{1}\tilde{\times}S^{1}\tilde{\times}I$ is homeomorphic to $M(0,1;1/2,1/2)$ and $M(-1,1;)$.
@@ 2.5. Hyperbolic 3-manifolds
A hyperbolic manifold is a complete Riemannian manifold of constant sectional curvature $-1$. We refer to [3, 55, 60] for details on hyperbolic 3-manifolds.
@@
An important property of a hyperbolic 3-manifold is the Mostow rigidity (also known as the Mostow–Prasad rigidity) [44, 54], which states that the hyperbolic structure on an atoroidal 3-manifold is unique. More precisely, suppose that $M_{0}$ and $M_{1}$ are finite-volume hyperbolic 3-manifolds. If there exists a homeomorphism $f\colon M_{0}\to M_{1}$, then $f$ is isotopic to a unique isometry from $M_{0}$ to $M_{1}$.
A link $L$ in a 3-manifold $X$ with (possibly empty) boundary consisting of tori is called hyperbolic if the interior of the complement $X\setminus\!\!\setminus L$ admits a finite-volume hyperbolic structure. Due to Thurston’s hyperbolization, plenty of links are hyperbolic. The hyperbolic structure of each hyperbolic link is unique by the Mostow rigidity. Hence the hyperbolic structure induces many geometric invariants, such as the hyperbolic volume. In this paper, we use the Mostow rigidity to consider symmetries of hyperbolic links.
@@ 2.6. JSJ decomposition
The JSJ decomposition is a standard decomposition of a compact irreducible 3-manifold along incompressible tori as follows. A properly embedded surface $S$ in a 3-manifold $M$ is 2-sided if its normal bundle is trivial. If $S$ and $M$ are orientable, then $S\subset M$ is always 2-sided. A 2-sided surface $S$ is compressible if there is a disk $D\subset M$ such that $D\cap S=\partial D$ and $\partial D$ does not bound a disk in $S$. Such a disk $D$ is called a compressing disk. In this case, we can compress the surface $S$ along $D$ to a new surface obtained by removing a tubular neighborhood of $\partial D$ from $S$ and capping it with two disks parallel to $D$. A 2-sided surface $S$ other than $S^{2}$ or $D^{2}$ is incompressible if it is not compressible. The loop theorem [25, Theorem 3.1] implies that $S\subset M$ is incompressible if and only if the induced homomorphism $\pi_{1}(S)\to\pi_{1}(M)$ is injective.
@@ 3.1. Links in the solid torus
In this subsection, we consider links in the solid torus $S^{1}\times D^{2}$. We first define elementary links in the solid torus, which are analogue to the torus links in the 3-sphere. Let $S^{1}=\mathbb{R}/\mathbb{Z}$ and $D^{2}=\{(x,y)\in\mathbb{R}^{2}\mid x^{2}+y^{2}\leq 1\}$. For $(p,q)\in\mathbb{Z}^{2}\setminus\{(0,0)\}$, define
[…]
@@
Let $d_{i}=\gcd(p_{i},q_{i})$, $p_{i}^{\prime}=p_{i}/d_{i}$, and $q_{i}^{\prime}=q_{i}/d_{i}$. Suppose that $l_{0}=0$. Since $p_{0}$ does not divide $q_{0}$, any component of $T_{0}(p_{0},q_{0})$ is not isotopic to $T_{1}(0,0)$. Hence $l_{1}=0$. Since the number of component of the link $T_{0}(p_{i},q_{i})$ is $d_{i}$, we have $d_{0}=d_{1}$. Since any component of $T_{0}(p_{i},q_{i})$ represents $p_{i}^{\prime}$ times of a generator of $\pi_{1}(S^{1}\times D^{2})$, we have $p_{0}^{\prime}=p_{1}^{\prime}$. Moreover, the knots $T_{0}(p_{0}^{\prime},q_{0}^{\prime})$ and $T_{0}(p_{1}^{\prime},q_{1}^{\prime})$ are isotopic. Since the standard embedding of the solid torus in $\mathbb{R}^{3}$ maps $T_{0}(p,q)$ to the $(p,q)$-torus link, the $(p_{0}^{\prime},q_{0}^{\prime})$-torus knot and the $(p_{1}^{\prime},q_{1}^{\prime})$-torus knot are isotopic. The classification of torus knots in [45, Theorem 7.4.3] implies that $q_{0}^{\prime}=q_{1}^{\prime}$. Hence $p_{0}=p_{1}$ and $q_{0}=q_{1}$.
@@
where $\gcd(0,0)=0$. The complement of a link in the solid torus is homeomorphic to the complement of a link in the 3-sphere one of whose component is the unknot. Moreover, Burde and Murasugi [7] classified the links in the 3-sphere whose complements are Seifert fibered (see also [6, Proposition 3.3]). This implies the following classification of Seifert fibered links.
@@
In the latter case, the other region bounded by $T$ is called a knotted hole ball.
@@
Since $\pi_{1}(S^{1}\times D^{2})\cong\mathbb{Z}$ does not have a subgroup isomorphic to $\pi_{1}(T)\cong\mathbb{Z}\times\mathbb{Z}$, any torus $T$ in $S^{1}\times D^{2}$ is compressible. A sphere $S$ is obtained by compressing $T$ along a compressing disk $D$. The sphere $S$ bounds a ball $B$ since the solid torus is irreducible. If $B$ and $D$ are disjoint, then $T$ bounds a solid torus. If $D$ is contained in $B$, then $T$ bounds a 3-manifold obtained by attaching a 1-handle to $\overline{S^{1}\times D^{2}\setminus B}$, which is homeomorphic to the connected sum of two solid tori. ∎
@@
We show that the outermost JSJ piece can be reembedded to a link complement in the same manner as [6, Proposition 2.2].
@@
Suppose that $M$ has a boundary component $T$ which bounds a knotted hole ball as in Lemma 3.3. Then $M$ is contained in the other region bounded by $T$, which is obtained by attaching a 1-handle to the complement of a 3-ball in $S^{1}\times D^{2}$. By unknotting this 1-handle, we can reembed $M$ in $S^{1}\times D^{2}$ so that $T$ bounds a solid torus as shown in Figure 4. Note that these attached 1-handles are disjoint by [6, Proposition 2.1]. After these operations for all such $T$, the manifold $M$ is reembedded in $S^{1}\times D^{2}$ so that $S^{1}\times D^{2}\setminus M$ is a tubular neighborhood of links in $S^{1}\times D^{2}$. ∎
Figure 4. Reembedding of a JSJ piece $M$.
The classification of the Seifert fibered submanifold of $S^{3}$ in [6, Proposition 3.2] implies the following classification.

### v2

@@ 2.1. Links in 3-manifolds, periodic tangles, and minimal motifs
In Example 2.6, the motif (e) is minimal. The hyperbolic structures of the complements of the motifs (a)-(e) are obtained by gluing regular ideal octahedra (see [10] for details). In particular, the complement of (e) can be decomposed into two regular ideal octahedra, and its volume is $2v_{\mathrm{oct}}$, where $v_{\mathrm{oct}}=3.66$... is the volume of a regular ideal octahedron. Since the complement of every motif of a hyperbolic doubly periodic tangle has at least three cusps, the volume is more than $v_{\mathrm{oct}}$ by [2]. Hence the complement of (e) cannot be a non-trivial finite cover of any other motif.
@@ 2.3. Prime decomposition
The connected sum $M_{0}\#M_{1}$ of 3-manifolds $M_{0}$ and $M_{1}$ is obtained by removing a ball in each $M_{i}$ and gluing them along the resulting boundary spheres. If $M_{0}$ and $M_{1}$ are connected and oriented, the oriented connected sum $M_{0}\#M_{1}$ is uniquely determined. The connected sum operation is commutative and associative, and $S^{3}$ is the identity element for the connected sum. A 3-manifold $M$ is prime if M=P#QM=P\#Q implies that $P=S^{3}$ or $Q=S^{3}$. A 3-manifold $M$ is irreducible if every embedded sphere in $M$ bounds a 3-ball. In the context of our paper, the 3-manifolds $S^{3}$, $S^{1}\times D^{2}$, $T^{2}\times I$, and $T^{3}$ are irreducible. If an orientable 3-manifold $M$ is prime and not irreducible, then $M=S^{1}\times S^{2}$. The prime decomposition of a compact orientable 3-manifold $M$ is a decomposition of the form $M=P_{1}\#\dots\#P_{n}$, where each $P_{i}$ is a prime 3-manifold. Moreover, a prime decomposition is unique up to permutation and connected sums with additional copies of $S^{3}$.
Let $L$ be a link in an irreducible 3-manifold $X$. We say that $L$ splits to links $L^{\prime}$, $L_{1}$, …, $L_{m}$ if $L=L^{\prime}\sqcup L_{1}\sqcup\dots\sqcup L_{m}$ and there are disjoint balls $B_{1},\dots,B_{m}$ in $X$ disjoint from $L^{\prime}$ such that each $B_{i}$ contains $L_{i}$. If $m\geq 1$, then $L$ is called a split link. The link $L$ is split if and only if the complement $X\setminus\!\!\setminus L$ is not irreducible. The following lemma is a consequence of the prime decomposition of the complement $X\setminus\!\!\setminus L$ of a link $L$.
@@ 2.4. Seifert fibered 3-manifolds
For coprime integers $p$ and $q\neq 0$, a $(p,q)$-model Seifert fibration on $S^{1}\times D^{2}$ is a decomposition of $S^{1}\times D^{2}$ into disjoint circles, called fibers, obtained from the segments $0,1]\times\{x\}$ in $0,1]\times D^{2}$ by gluing $\{0\}\times D^{2}$ and $\{1\}\times D^{2}$ via a rotation of $2\pi p/q$. The $(p,1)$-model Seifert fibration is the same as the fibration given by the usual product structure on $S^{1}\times D^{2}$ (sometimes referred to as a product fibration) – that is, every fiber is a circle over a point in the disk and the resulting space is $S^{1}\times D^{2}$.
@@
The classification of Seifert fibered 3-manifolds was established many years ago, and here, we use the notation for orientable Seifert fibered 3-manifolds following [7, 29]. For $g\in\mathbb{Z}$ and $b\in\mathbb{Z}_{\geq 0}$, let $\Sigma_{g,b}$ denote the orientable compact surface of genus $g$ with $b$ boundary components if $g\geq 0$. Let $\Sigma_{g,b}$ denote the non-orientable compact surface that is the connected sum of g-g copies of $\mathbb{R}P^{2}$ and $\Sigma_{0,b}$ if $g<0$. Let $M(g,b;\alpha_{1}/\beta_{1},\dots,\alpha_{k}/\beta_{k})$ denote the orientable Seifert fibered 3-manifold that is fibered over $\Sigma_{g,b}$ with at most $k$ singular fibers and fiber data $\alpha_{i}/\beta_{i}\in\mathbb{Q}$. More precisely, the 3-manifold $M(g,b;\alpha_{1}/\beta_{1},\dots,\alpha_{k}/\beta_{k})$ is obtained from the orientable $S^{1}$-bundle over $\Sigma_{g,b+k}$ by Dehn filling along $k$ boundary components so that the meridians of attached solid tori are glued to the slopes $\alpha_{i}/\beta_{i}$. Here the slope $\alpha_{i}/\beta_{i}$ on the torus $T_{i}$ for coprime integers $\alpha_{i}$ and $\beta_{i}$ is represented by $\alpha_{i}\lambda_{i}+\beta_{i}\mu_{i}\in H_{1}(T_{i},\mathbb{Z})$ for an oriented pair of a fiber $\lambda_{i}$ and a boundary component of a section $\mu_{i}$. The Seifert fibration on $M(g,b;\alpha_{1}/\beta_{1},\dots,\alpha_{k}/\beta_{k})$ is obtained by extending the product fibration over $\Sigma_{g,b+k}$. The 3-manifolds with non-unique Seifert fibrations are classified. We state the classification for bounded orientable Seifert fibered 3-manifolds.
@@
The solid torus $S^{1}\times D^{2}$ admits the model Seifert fibrations of $M(0,1;\alpha/\beta)$. The thickened torus $T^{2}\times I$ admits the product fibration of $M(0,2;)$, which is unique up to isomorphism. The twisted $I$-bundle over the Klein bottle $S^{1}\tilde{\times}S^{1}\tilde{\times}I$ is homeomorphic to $M(0,1;1/2,1/2)$ and M(1,1;)M(-1,1;).
@@ 2.5. Hyperbolic 3-manifolds
A hyperbolic manifold is a complete Riemannian manifold of constant sectional curvature 1-1. We refer to [4, 64, 69] for details on hyperbolic 3-manifolds.
@@
An important property of a hyperbolic 3-manifold is the Mostow rigidity theorem (also known as the Mostow–Prasad rigidity theorem) [53, 63], which states that the hyperbolic structure on an atoroidal 3-manifold is unique. More precisely, suppose that $M_{0}$ and $M_{1}$ are finite-volume hyperbolic 3-manifolds. If there exists a homeomorphism $f\colon M_{0}\to M_{1}$, then $f$ is isotopic to a unique isometry from $M_{0}$ to $M_{1}$. In fact, a homotopy equivalence between finite-volume hyperbolic manifolds of dimension at least three is homotopic to an isometry. The fact that homotopic homeomorphisms between Haken 3-manifolds are isotopic was shown by Waldhausen [71] (and shown by Gabai, Meyerhoff, and Thurston [19] for general 3-manifolds).
A link $L$ in a 3-manifold $X$ with (possibly empty) boundary consisting of tori is called hyperbolic if the interior of the complement $X\setminus\!\!\setminus L$ admits a finite-volume hyperbolic structure. Due to Thurston’s hyperbolization, plenty of links are hyperbolic. The hyperbolic structure of each hyperbolic link is unique by the Mostow rigidity theorem. Hence the hyperbolic structure induces many geometric invariants, such as the hyperbolic volume. In this paper, we use the Mostow rigidity theorem to consider symmetries of hyperbolic links.
@@ 2.6. JSJ decomposition
The JSJ decomposition is a standard decomposition of a compact irreducible 3-manifold along incompressible tori as follows. A properly embedded surface $S$ in a 3-manifold $M$ is 2-sided if its normal bundle is trivial. If $S$ and $M$ are orientable, then $S\subset M$ is always 2-sided. A 2-sided surface $S$ is compressible if there is a disk $D\subset M$ such that $D\cap S=\partial D$ and $\partial D$ does not bound a disk in $S$. Such a disk $D$ is called a compressing disk. In this case, we can compress the surface $S$ along $D$ to a new surface obtained by removing a tubular neighborhood of $\partial D$ from $S$ and capping it with two disks parallel to $D$. A 2-sided surface $S$ other than $S^{2}$ or $D^{2}$ is incompressible if it is not compressible. The loop theorem [29, Theorem 3.1] implies that $S\subset M$ is incompressible if and only if the induced homomorphism $\pi_{1}(S)\to\pi_{1}(M)$ is injective.
@@
To consider the JSJ decomposition of a link complement $X\setminus\!\!\setminus L$, we need to describe regions bounded by an embedded torus in $X$. We say that a knotted hole ball in a 3-manifold $X$ is the complement of a properly embedded arc in a 3-ball contained in $X$ that is not a solid torus.
@@ 3.1. Links in the solid torus
In this subsection, we consider links in the solid torus $S^{1}\times D^{2}$. We first define elementary links in the solid torus, which are analogous to the torus links in the 3-sphere. Let $S^{1}=\mathbb{R}/\mathbb{Z}$ and $D^{2}=\{(x,y)\in\mathbb{R}^{2}\mid x^{2}+y^{2}\leq 1\}$. For $(p,q)\in\mathbb{Z}^{2}\setminus\{(0,0)\}$, define
$T_{0}(p,q)=\{$ $\left(p^{\prime}t,\frac{1}{2}\cos 2\pi\left(q^{\prime}t+\frac{k}{p}\right),\frac{1}{2}\sin 2\pi\left(q^{\prime}t+\frac{k}{p}\right)\right)\in S^{1}\times D^{2}\mid$
@@
Let $d_{i}=\gcd(p_{i},q_{i})$, $p_{i}^{\prime}=p_{i}/d_{i}$, and $q_{i}^{\prime}=q_{i}/d_{i}$. Suppose that $l_{0}=0$. Since $p_{0}$ does not divide $q_{0}$, any component of $T_{0}(p_{0},q_{0})$ is not isotopic to $T_{1}(0,0)$. Hence $l_{1}=0$. Since the number of component of the link $T_{0}(p_{i},q_{i})$ is $d_{i}$, we have $d_{0}=d_{1}$. Since any component of $T_{0}(p_{i},q_{i})$ represents $p_{i}^{\prime}$ times of a generator of $\pi_{1}(S^{1}\times D^{2})$, we have $p_{0}^{\prime}=p_{1}^{\prime}$. Moreover, the knots $T_{0}(p_{0}^{\prime},q_{0}^{\prime})$ and $T_{0}(p_{1}^{\prime},q_{1}^{\prime})$ are isotopic. Since the standard embedding of the solid torus in $\mathbb{R}^{3}$ maps $T_{0}(p,q)$ to the $(p,q)$-torus link, the $(p_{0}^{\prime},q_{0}^{\prime})$-torus knot and the $(p_{1}^{\prime},q_{1}^{\prime})$-torus knot are isotopic. The classification of torus knots in [54, Theorem 7.4.3] implies that $q_{0}^{\prime}=q_{1}^{\prime}$. Hence $p_{0}=p_{1}$ and $q_{0}=q_{1}$.
@@
where $\gcd(0,0)=0$. The complement of a link in the solid torus is homeomorphic to the complement of a link in the 3-sphere one of whose components is the unknot. Moreover, Burde and Murasugi [8] classified the links in the 3-sphere whose complements are Seifert fibered (see also [7, Proposition 3.3]). This implies the following classification of Seifert fibered links.
@@
In the latter case, $T$ bounds a solid torus or a knotted hole ball.
@@
Since $\pi_{1}(S^{1}\times D^{2})\cong\mathbb{Z}$ does not have a subgroup isomorphic to $\pi_{1}(T)\cong\mathbb{Z}\times\mathbb{Z}$, any torus $T$ in $S^{1}\times D^{2}$ is compressible. A sphere $S$ is obtained by compressing $T$ along a compressing disk $D$. The sphere $S$ bounds a ball $B$ since the solid torus is irreducible. If $B$ and $D$ are disjoint, then $T$ bounds a solid torus. If $D$ is contained in $B$, then $T$ bounds a 3-manifold obtained by attaching a 1-handle to $\overline{S^{1}\times D^{2}\setminus B}$, which is homeomorphic to the connected sum of two solid tori. In this case, the other region bounded by $T$ is the complement of the 1-handle in $B$, and so it is a solid torus or a knotted hole ball. ∎
@@
We show that the outermost JSJ piece can be re-embedded as a link complement in the same manner as [7, Proposition 2.2].
@@
Suppose that $M$ has a boundary component $T$ which bounds a knotted hole ball as in Lemma 3.3. Then $M$ is contained in the other region bounded by $T$, which is obtained by attaching a 1-handle to the complement of a 3-ball in $S^{1}\times D^{2}$. By unknotting this 1-handle, we can re-embed $M$ in $S^{1}\times D^{2}$ so that $T$ bounds a solid torus as shown in Figure 4. Note that these attached 1-handles are disjoint by [7, Proposition 2.1]. After these operations for all such $T$, the manifold $M$ is re-embedded in $S^{1}\times D^{2}$ so that $S^{1}\times D^{2}\setminus M$ is a tubular neighborhood of links in $S^{1}\times D^{2}$. ∎
Figure 4. Re-embedding of a JSJ piece $M$.
The classification of the Seifert fibered submanifold of $S^{3}$ in [7, Proposition 3.2] implies the following classification.